If Sin x + cos x = a then | sin x - cos x | =1) √1-a square2) √a square -13) √2-a square4) √a square - 2

Grade:8

2 Answers

manna
7 years ago
sinx+cosx=a
take a sqr both side
(sinx+cosx)^2=a^2
sin^2x+cos^2x+2sinx*cosx=a^2------------------(a+b)^2=a^2+b^2+2ab
1+2sinx*cosx=a^2--------------------------sin^2x+cos^2x=1
sin2x=a^2-1
now i would encourage you to do it yourself onwords.....
show your work....
aadi
6 years ago
sin^2x+cos^2x+2sinxcosx=a^2
sin^2x+cos^2x=a^2-2sinxcosx
-2sinxcosx=sin^2x+cos^2x-a^2
(sinx-cosx)^2=2(sin^2x+cos^2x)-a^2
sinx-cosx=sqrt(2-a^2)